# ISAC channel simulator GBSM

### v3

@@ 2.2.1 RCS polarization
where and denote the horizontally and vertically polarized components of the received electric field, respectively, and denote the horizontally and vertically polarized components of the transmitted electric field, respectively. The scattering matrix S represents the coefficients related to the target’s scattering characteristics and polarization.
@@ 2.2.2 RCS modeling based Measurement
where component is included in the large scale and is a deterministic value. The component is included in the small scale and is a deterministic angle-dependent value or scattering pattern based on measurements. The component is a normal distribution with a mean and variance. Modeling the RCS of a UAV based on this framework, since its RCS is angle-dependent, the component is set to the mean or peak value, or simply set to 1, etc., the component is a function related to the angle, and is a random value independent of the angle, representing the fluctuation of RCS. This model framework can effectively model the RCS of sensing targets with different scattering patterns, so we simulate with subsequent discussions on RCS modeling for different sensing targets based on this model framework.
Different from the human body and small UAV, vehicle RCS and large UAV exhibit spiking characteristics that render them unsuitable for modeling with omnidirectional patterns. The RCS of humans and small UAVs can be modeled using and with a log-normal distribution, in which the mean and variance are determined from empirical measurements. The RCS of vehicles and large UAVs displays four distinct peaks in its angular distribution, highlighting the need for further investigation into and simulation methodologies. The conversion formula for single-station and dual-station RCS values is as follows:
This equation is valid under the condition that the target is sufficiently smooth and free from shadowing and back-scatter effects. It necessitates validation through measurement results and can be modeled analogous to an antenna radiation pattern.
@@ 3.2 Detection metric
Substituting equation (23) into equation (22) results in
Further, the relationship between and is obtained as The probability of detection, , represents the likelihood of correctly detecting a target when one is present and is defined as follows
@@ 4 Simulation Results
In the simulation experiments, the primary focus is on the reliability of the ISAC channel cascading method and the accuracy of the RCS simulation results. RAN1 118 bis meeting has agreed 8 options for concatenating the Tx-target and target-Rx link in the target channel as
@@
Option 2A: ray level 1-by-1 coupling between Tx-target and target-Rx link for Case 1/2/3/4.
@@
| CaseA | Concatenation only for radio propagation NLOS-NLOS |
@@
Options 0, 1, 2, and 3A in #118-bis simulation options for concatenation are uniformly represented by CaseA. The ray-level 1-by-1 coupling in order is excluded, as it is equivalent to Case 2O. Additionally, the specific coupling method of ray 1-by-1, whether sequential or random, does not affect Case 1, Case 2O, or Case 2R. This is because the power of each ray pair depends solely on the power of the corresponding clusters, with rays within the same cluster having equal power. Therefore, the coupling method at the ray level does not affect the distribution, as the power of each ray pair remains the same once the cluster-level concatenation is complete. The simulation configuration parameters and corresponding results are presented as Table 4.
@@
Figure 4 illustrates the human target’s RCS simulation and measurement results at 5 GHz, 14 GHz, and 28 GHz in Mono-static mode. The simulated results demonstrate a good alignment with the measured results reported in reference [25], thereby validating the feasibility of the log-normal fitting scheme. Additionally, this finding confirms the differences in RCS distribution between ISAC channels and traditional radar systems. This discrepancy arises from the fact that traditional radar RCS perception typically occurs at greater distances, while ISAC services are concentrated on closer ranges. Consequently, it results in variations between the RCS distributions observed in radar applications and those measured in ISAC.
@@ 5 Conclusion
Through down-selection simulations of the 8 concatenation options provided in the #118-bis meeting, we observed the relation between different cascading methods compared to the fully convoluted baseline. NLOS-NLOS power normalization should be performed, as it ensures overall small-scale power normalization and brings the statistical characteristics of different cascading options closer to option 0. After NLOS-NLOS power normalization, the statistical distribution of option 1 matches exactly with that of option 0. In option 2, the random matching method performs better than the sequential coupling method. Option 3 fits the fully convoluted result better than option 2. However, option 3 abandons the traditional cluster concept. If the cluster concept is necessary (e.g., for MIMO OTA testing), we recommend using option 2, which retains the cluster concept while closely approximating the fully convoluted results.

### v4

@@ 2.2.1 RCS polarization
The scattering matrix S represents the coefficients related to the target’s scattering characteristics and polarization as follows:
where, the initial random phase is uniformly distributed within . For , there are three primary perspectives within RAN1: , , and . Here, , , , and are variables generated for each path, and their specific values remain subject to further study.
@@ 2.2.2 RCS modeling based Measurement
where component is included in the large scale and is a deterministic value. The component is included in the small scale and is a deterministic angle-dependent value or scattering pattern based on measurements. The component is a normal distribution with a mean and variance. is used as in (1), while and are used as in (2.1.2), (2.1.2), (2.1.2), (6). Modeling the RCS of a UAV based on this framework, since its RCS is angle-dependent, the component is set to the mean or peak value, or simply set to 1, etc., the component is a function related to the angle, and is a random value independent of the angle, representing the fluctuation of RCS. This model framework can effectively model the RCS of sensing targets with different scattering patterns. In section 4, we present the measured results of the human target RCS and the simulation fitting results based on the proposed framework, which are used to comparatively validate the feasibility of the model.
Different from the human body and small UAV, vehicle RCS and large UAV exhibit spiking characteristics that render them unsuitable for modeling with omnidirectional patterns. The RCS of humans and small UAVs can be modeled using and with a log-normal distribution, in which the mean and variance are determined from empirical measurements. The RCS of vehicles and large UAVs displays four distinct peaks in its angular distribution, highlighting the need for further investigation into and simulation methodologies.
The investigation of both Mono-static and Bi-static RCS is of equal importance, and their relationship warrants further discussion in RAN1. As illustrated in Figure 3, Mono-static sensing and Bi-static sensing are presented. Notably, Mono-static sensing can be regarded as a special case of Bi-static sensing when . The equivalence relationship between Mono-static and Bi-static RCS was initially explored in radar research. Kell proposed a general scattering center-based Mono-static-to-Bi-static equivalence theorem (MBET) [28]. This theorem demonstrates that the Bi-static RCS can be approximated by the Mono-static RCS measured at a frequency scaled by on the bisector of the Bi-static angle , as expressed by . where is the receiver orientation angle, and is the carrier frequency. When the Bi-static angle is small, the Bi-static RCS can be approximated by the Mono-static RCS as follows:
According to a comprehensive investigation [29], the MBET demonstrates strong performance for simple geometries, maintaining accuracy for Bi-static angles of up to 30° (including sidelobe structure and amplitudes). However, for minimally complex objects, the accuracy of the MBET diminishes significantly, remaining reliable only for Bi-static angles less than 15–20°. Future research is needed to develop modeling methods for Bi-static RCS at larger Bi-static angle .
@@ 3.2 Detection metric
Here, represents the standard deviation of the noise. Substituting equation (24) into equation (23) results in
Further, the relationship between and is obtained as . The probability of detection represents the likelihood of correctly detecting a target when one is present and is defined as follows
Here, denotes the signal amplitude, and represents the modified Bessel function of the first kind. Further exploration of the relationship between the proposed sensing performance metrics and traditional system performance metrics, such as bit error rate (BER), symbol error rate (SER), and signal-to-noise ratio (SNR), holds significant reference value for ISAC system design and testing. Considering that both SER and BER are also related to SNR, the aforementioned accuracy metrics are directly influenced by the parameters used in the detection process such as ranging and delay, velocity and Doppler, and angle and phase. Therefore, we will focus on investigating the interplay between SNR and detection metrics (probability of detection and false alarm probability).
As shown in Figure 4(a), the false alarm probability decreases exponentially with respect to the threshold level. When the noise distribution function is fixed, the magnitude of the false alarm probability is entirely dependent on the threshold level, which decreases exponentially as the threshold increases. As shown in Figure 4(b), for a fixed false alarm probability, the detection probability increases with increasing SNR, i.e., for a constant threshold level, the detection probability increases as the SNR increases. In other words, for a fixed SNR, a lower false alarm probability corresponds to a lower detection probability, and a higher false alarm probability corresponds to a higher detection probability. By combining Figure 4(a) and 4(b), it is evident that when the relative threshold increases, the false alarm probability decreases, but the detection probability also decreases. We always aim to improve the detection probability for a fixed false alarm probability, which can only be achieved by increasing the SNR.
@@ 4 Simulation Results
4.1 Concatenation methods
In the simulation experiments, the primary focus is on the reliability of the ISAC channel concatenated method and the accuracy of the RCS simulation results. RAN1 #118bis meeting has agreed 8 options for concatenating the Tx-target and target-Rx link in the target channel as
@@ 4.1 Concatenation methods
Option 3: ray level 1-by-1 coupling between Tx-target and target-Rx link for Case 1/2/3/4.
@@
CaseA | Concatenation only for NLOS+NLOS propagation condition |
@@
Options 0, 1, 2, and 3A in #118-bis simulation options for concatenation are uniformly represented by CaseA. The ray-level 1-by-1 coupling in order is excluded, as it is equivalent to Case 2O. Additionally, the specific coupling method of ray 1-by-1, whether sequential or random, does not affect Case 1, Case 2O, or Case 2R. This is because the power of each ray pair depends solely on the power of the corresponding clusters, with rays within the same cluster having equal power. Therefore, the coupling method at the ray level does not affect the distribution, as the power of each ray pair remains the same once the cluster-level concatenation is complete. The simulation configuration parameters and corresponding results are presented in Table 4.
@@
4.2 RCS measurement and simulation
We have planned a series of RCS measurement campaigns to acquire data for supporting the precise modeling of target RCS. Vehicles, humans, and UAVs have been selected as the sensing targets for measurement, and these activities will be conducted in an anechoic chamber. An anechoic chamber is utilized to effectively minimize environmental interference. Within this controlled environment, RCS measurements are conducted for various sensing targets, including UAVs, humans, and vehicles. The measurement system comprises two horn antennas, which function as the sensing transmitter and receiver, a Vector Network Analyzer (VNA), a power amplifier, a PC controller, and other components, as illustrated in Figure 6. The two horn antennas are positioned in close proximity (with a spacing of approximately 10 cm) to emulate the Mono-static sensing mode, while the sensing target is placed at the center of an automatic turntable. By controlling the rotation of the turntable, the target’s echo is measured at different angles. The turntable rotates 360° clockwise in 5° increments. Detailed measurement parameters are provided in Table 5.
Parameters | Value/Type |
Carrier frequency | 36GHz, 28GHz, 20GHz, 15GHz, 10GHz |
Bandwidth | 3GHz |
Tx/Rx antenna type | Horn/Horn |
Polarization | Vertical |
Tx/Rx distance form target | 6m, 12m |
Sensing mode | Mono-static |
Target | Human, Vehicle, UAV |
Based on the RCS measurement results, Figure 7 presents the cumulative distribution functions (CDFs) of the RCS for humans, vehicles, and UAVs at different frequencies, respectively. Each CDF is fitted using a normal distribution function. During this process, angles exhibiting strong directionality for vehicles and UAVs are excluded, while angles with similar RCS values are retained. Subsequently, a normal distribution function is applied to fit the data, enabling an investigation into the frequency dependence of the sensing targets. In Figure 7, it can be observed that the RCS of the three sensing targets—humans, vehicles, and UAVs—exhibits a certain frequency-dependent characteristic, where the RCS values increase with rising frequency. For the vehicle target, however, there is a noticeable deviation between the fitted normal distribution and the measured values. This discrepancy arises because the vehicle’s relatively large size and the significant differences in the surface areas of its four sides introduce a deterministic angle-dependent component to its RCS. Further research is required to explore modeling methods for the deterministic component of the vehicle’s RCS. In contrast, the RCS values of human and UAV targets can be well-fitted using a log-normal distribution. Based on the measurement results, we provide reference tables for the fitted and values of humans and UAVs under various conditions. Due to the extensive data, specific values can be obtained from proposals [26] and [27].
@@ 5 Conclusion
Through down-selection simulations of the 8 concatenation options provided in the #118bis meeting, we observed the relation between different cascading methods compared to the fully convoluted baseline. NLOS-NLOS power normalization should be performed, as it ensures overall small-scale power normalization and brings the statistical characteristics of different cascading options closer to option 0. After NLOS-NLOS power normalization, the statistical distribution of option 1 matches exactly with that of option 0. In option 2, the random matching method performs better than the sequential coupling method. Option 3 demonstrates excellent performance in fitting the fully convoluted result. However, option 3 abandons the traditional cluster concept. If the cluster concept is necessary (e.g., for MIMO OTA testing), we recommend using option 2, which retains the cluster concept while closely approximating the fully convoluted results.